# Generative AI and Large Language Models for Cyber Security: All Insights You Need

### v1

@@ Abstract
The rapid evolution of cyber threats requires innovative approaches to enhance cybersecurity defenses. In this paper, we provide a comprehensive and in-depth review of the future of cybersecurity through the lens of Generative AI and Large Language Models (LLMs). We explore LLM applications across various cybersecurity domains, including hardware design security, intrusion detection, software engineering, design verification, cyber threat intelligence, malware detection, and phishing and spam detection. We present a detailed overview of LLM evolution and its current state, focusing on advancements in 42 specific models, such as GPT-4o, GPT-4, GPT-3.5, Mixtral-8x7B, BERT, Falcon2, Gemma, Phi-2, Phi-3, and LLaMA. Our analysis extends to the vulnerabilities inherent in LLMs, including prompt injection, insecure output handling, training and inference data poisoning, DDoS attacks, and adversarial natural language instructions. We also delve into mitigation strategies to protect these models, providing a comprehensive look at potential attack scenarios and prevention techniques. Furthermore, we evaluate the performance of 42 LLM models in cybersecurity knowledge and hardware security, highlighting their strengths and weaknesses. Our study thoroughly evaluates cybersecurity datasets tailored for LLM training and testing, covering the entire lifecycle from data creation to usage and identifying gaps and opportunities for future research. We discuss the challenges and limitations of employing LLMs in cybersecurity settings, such as dealing with adversarial attacks and ensuring model robustness. In addition, we review the new strategies for leveraging LLMs, including advanced techniques like Half-Quadratic Quantization (HQQ), Reinforcement Learning with Human Feedback (RLHF), Direct Preference Optimization (DPO), Odds Ratio Preference Optimization (ORPO), GPT-Generated Unified Format (GGUF), Quantized Low-Rank Adapters (QLoRA), and Retrieval-Augmented Generation (RAG). These insights aim to enhance real-time cybersecurity defenses and improve the sophistication of LLM applications in threat detection and response. Our paper aims to provide a foundational understanding and strategic direction for integrating LLMs into future cybersecurity frameworks, emphasizing the importance of innovation and robust model deployment to safeguard against evolving cyber threats.
@@ I Introduction
The history of Natural Language Processing (NLP) dates back to the 1950s when the Turing test was developed. However, NLP has seen significant advancements in recent decades with the introduction of Recurrent Neural Networks (RNN) [1], Long Short-Term Memory (LSTM) [2], Gated Recurrent Units (GRU) [3], and Transformer methods [4]. RNN was first introduced in the 1990s to model data sequences. LSTM, a variant of RNN, was introduced in 1997, which addressed the vanishing gradient problem and allowed for longer-term memory in NLP models. GRU, another variant of RNN, was introduced in 2014, which reduced the number of parameters and improved computational efficiency [5]. The latest breakthrough in NLP was the introduction of Transformers in 2017, enabling parallel processing of sequential data and revolutionizing tasks like machine translation. These methods have significantly improved various NLP tasks, including sentiment analysis, language generation, and translation [4, 6, 7].
@@
Cybersecurity professionals often need to sift through a vast amount of textual data, including security alerts, incident reports, threat feeds, and research papers, to stay ahead of evolving threats. LLMs, like Falcon 180b [16], possess natural language understanding capabilities that enable them to parse, summarize, and contextualize this information efficiently [17, 18, 7]. They can assist in rapidly identifying relevant threat intelligence, allowing analysts to make more informed decisions and prioritize responses [19]. LLMs can excel in various domains within cybersecurity. Figure 1 highlights the top nine use cases and applications for LLMs in this field [20].
@@ II Related reviews
This section delves into a curated collection of recent articles that significantly contribute to the evolving landscape of LLMs and their multifaceted applications. These reviews offer a comprehensive and insightful exploration into various dimensions of LLMs, including their innovative applications in hardware design security, evaluation methodologies, and evolving role in artificial intelligence. Further, they cover cutting-edge advancements in Pre-trained Language Models (PLMs) for NLP, delve into the intricacies of instruction fine-tuning for LLMs, and explore their impactful integration into software engineering. The section also encompasses an in-depth look at multimodal algorithms, examines the critical aspect of alignment requirements for LLMs, and discusses integrating external knowledge into PLMs to enhance NLP tasks. Lastly, it sheds light on the burgeoning field of Controllable Text Generation (CTG) in Natural Language Generation (NLG), highlighting the latest trends and challenges in this dynamic and rapidly advancing area of research [41, 42, 43].
II-A Applications of LLMs in Hardware Design Security
II-B Evaluation of LLMs
@@
II-C Evolution and State of LLMs in AI
@@ II-D Advancements in PLMs for NLP
II-E Instruction Fine-Tuning for LLMs
@@
II-F LLMs in Software Engineering
@@ II-G Multimodal Algorithms
II-H Alignment Requirements for LLMs
@@
II-I Knowledge-Enhanced Pre-trained Language Models
@@
II-J Controllable Text Generation in NLG
@@ II-K LLM for Cyber Security
II-L Our survey compared to related surveys
@@ III-B Transformer models
TABLE II: LLMs-based models for Cyber Security (Part I).
@@
Ferrag and Leandros [59] 2019 DeepCoin (Blockchain and Deep Learning) CICID2017 dataset, Power system dataset, Bot-IoT dataset Energy framework for Smart Grids Proposal of DeepCoin framework combining blockchain and deep learning for smart grid security The paper does not address the potential scalability issues that may arise as the number of nodes in the network increases
@@
TABLE III: LLMs-based models for Cyber Security (Part II).
@@ IV LLMs-based models for Cyber Security
This section reviews recent studies employing LLM-based models (i.e., Recurrent Neural Networks-based and transformer-based models) for threat detection, malware classification, intrusion detection, and software vulnerability detection. Tables II, III, and Figure 4 present the LLM-based solutions for Cyber Security Use Cases.
@@ IV-B3 Intrusion Detection
Wu et al. [53] proposed an RTID that reconstructs feature representations in imbalanced datasets to make a trade-off between dimensionality reduction and feature retention. The proposed method utilizes a stacked encoder-decoder neural network and a self-attention mechanism for network traffic type classification. The results with CICID2017 and CIC-DDoS2019 datasets demonstrate the proposed method’s effectiveness in intrusion detection compared to classical machine learning and deep learning algorithms. Ghourbi et al. [60] propose an intrusion and malware detection system to secure the entire network of the healthcare system independently of the installed devices and computers. The proposed solution includes two components: an intrusion detection system for medical devices installed in the healthcare network and a malware detection system for data servers and medical staff computers. The proposed system is based on optimized LightGBM and Transformer-based models. It is trained with four different datasets to ensure a varied knowledge of the different attacks affecting the healthcare sector. The experimental evaluation of the approach showed remarkable accuracies of 99%.
@@ IV-B7 Hardware Security Evaluation
Madhav et al. [86] evaluated the security of the HDL code generated by ChatGPT. The authors introduced a similar taxonomy to the NIST CWE . The authors conducted various experiments to explore the impact of prompt engineering on the security of the generated hardware design.
@@ IV-B9 Protocols verification
TABLE IV: Comparison of 19 LLMs Models’ Performance in Hardware Security Knowledge.
@@
TABLE V: Comparison of 42 LLMs Models’ Performance in Cyber Security Knowledge.

### v2

# Generative AI in Cybersecurity: A Comprehensive Review of LLM Applications and Vulnerabilities

@@ Abstract
This paper provides a comprehensive review of the future of cybersecurity through Generative AI and Large Language Models (LLMs). We explore LLM applications across various domains, including hardware design security, intrusion detection, software engineering, design verification, cyber threat intelligence, malware detection, and phishing detection. We present an overview of LLM evolution and its current state, focusing on advancements in models such as GPT-4, GPT-3.5, Mixtral-8x7B, BERT, Falcon2, and LLaMA. Our analysis extends to LLM vulnerabilities, such as prompt injection, insecure output handling, data poisoning, DDoS attacks, and adversarial instructions. We delve into mitigation strategies to protect these models, providing a comprehensive look at potential attack scenarios and prevention techniques. Furthermore, we evaluate the performance of 42 LLM models in cybersecurity knowledge and hardware security, highlighting their strengths and weaknesses. We thoroughly evaluate cybersecurity datasets for LLM training and testing, covering the lifecycle from data creation to usage and identifying gaps for future research. In addition, we review new strategies for leveraging LLMs, including techniques like Half-Quadratic Quantization (HQQ), Reinforcement Learning with Human Feedback (RLHF), Direct Preference Optimization (DPO), Quantized Low-Rank Adapters (QLoRA), and Retrieval-Augmented Generation (RAG). These insights aim to enhance real-time cybersecurity defenses and improve the sophistication of LLM applications in threat detection and response. Our paper provides a foundational understanding and strategic direction for integrating LLMs into future cybersecurity frameworks, emphasizing innovation and robust model deployment to safeguard against evolving cyber threats.
@@ I Introduction
The history of Natural Language Processing (NLP) dates back to the 1950s when the Turing test was developed. However, NLP has seen significant advancements in recent decades with the introduction of Recurrent Neural Networks (RNN) [1], Long Short-Term Memory (LSTM) [2], Gated Recurrent Units (GRU) [3], and Transformer methods [4]. RNN was first introduced in the 1990s to model data sequences. LSTM, a variant of RNN, was introduced in 1997, which addressed the vanishing gradient problem and allowed for longer-term memory in NLP models. GRU, another variant of RNN, was introduced in 2014, which reduced the number of parameters and improved computational efficiency [5]. The latest breakthrough in NLP was the introduction of Transformers in 2017, enabling parallel processing of sequential data and revolutionizing tasks like machine translation. These methods have significantly improved various NLP tasks, including sentiment analysis, language generation, and translation [4, 6, 7, 8].
@@
Cybersecurity professionals often need to sift through a vast amount of textual data, including security alerts, incident reports, threat feeds, and research papers, to stay ahead of evolving threats. LLMs, like Falcon 180b [17], possess natural language understanding capabilities that enable them to parse, summarize, and contextualize this information efficiently [18, 19, 7]. They can assist in rapidly identifying relevant threat intelligence, allowing analysts to make more informed decisions and prioritize responses [20]. LLMs can excel in various domains within cybersecurity [21, 22]. Figure 1 highlights the top nine use cases and applications for LLMs in this field [23, 24].
@@ II Related reviews
This section delves into a curated collection of recent articles that significantly contribute to the evolving landscape of LLMs and their multifaceted applications. These reviews offer a comprehensive and insightful exploration into various dimensions of LLMs, including their innovative applications in hardware design security, evaluation methodologies, and evolving role in artificial intelligence. Further, they cover cutting-edge advancements in Pre-trained Language Models (PLMs) for NLP, delve into the intricacies of instruction fine-tuning for LLMs, and explore their impactful integration into software engineering. The section also encompasses an in-depth look at multimodal algorithms, examines the critical aspect of alignment requirements for LLMs, and discusses integrating external knowledge into PLMs to enhance NLP tasks. Lastly, it sheds light on the burgeoning field of Controllable Text Generation (CTG) in Natural Language Generation (NLG), highlighting the latest trends and challenges in this dynamic and rapidly advancing area of research [45, 46, 47]. Table I presents a comprehensive summary of existing reviews on LLMS across various application domains.
II-A Evaluation of LLMs
@@
II-B Evolution and State of LLMs in AI
@@ II-C Advancements in PLMs for NLP
II-D Instruction Fine-Tuning for LLMs
@@
II-E LLMs in Software Engineering
@@ II-F Multimodal Algorithms
II-G Alignment Requirements for LLMs
@@
II-H Knowledge-Enhanced Pre-trained Language Models
@@
II-I Controllable Text Generation in NLG
@@ II-J LLM for Cyber Security
II-K Our survey compared to related surveys
@@ III-B Transformer models
TABLE II: RNN-based models for Cyber Security.
Ferrag and Leandros [59] 2019 Blockchain and RNN CICID2017 dataset, Power system dataset, Bot-IoT dataset Energy framework for Smart Grids Proposal of DeepCoin framework combining blockchain and deep learning for smart grid security The paper does not address the potential scalability issues that may arise as the number of nodes in the network increases
TABLE III: Transformer-based models for Cyber Security (Part I).
@@
TABLE IV: Transformer-based models for Cyber Security (Part II).
@@
Patsakis et al. [84] 2024 LLM fine-tuned for deobfuscation tasks Malicious scripts from the Emotet malware campaign Malware Classification Demonstrated 69.56% accuracy in extracting URLs and 88.78% for domains of droppers; explored LLM potential in malware deobfuscation and reverse engineering Optimizing LLM fine-tuning for improved accuracy and integrating deobfuscation capabilities into operational security pipelines
Guo et al. [85] 2024 Fine-tuned open-source and general-purpose LLMs for binary classification Compiled dataset and five benchmark datasets for vulnerability detection Software Security Demonstrated fine-tuning’s effectiveness in improving detection accuracy; highlighted limitations of existing benchmark datasets Addressing dataset mislabeling and improving generalizability of models to unseen code scenarios
@@
Lykousas and Patsakis [86] 2024 LLMs for detecting hard-coded credentials in source code Public code repositories with embedded secrets and passwords Authentication and Code Security Highlighted differences in password patterns between developers and users; evaluated LLMs for detecting hard-coded credentials and discussed their limitations Improving LLM accuracy in detecting secrets and addressing context-sensitive password vulnerabilities
Karlsen et al. [87] 2024 Fine-tuned LLMs for sequence classification (e.g., DistilRoBERTa, GPT-2, GPT-Neo) Six datasets from web application and system logs Cybersecurity Log Analysis Proposed a new pipeline leveraging 60 fine-tuned models for log analysis; DistilRoBERTa achieved an F1-score of 0.998, outperforming state-of-the-art techniques Scaling models for more diverse log formats and optimizing for real-time analysis in dynamic environments
Mechri et al. [88] 2025 Decoder-only Transformer with 64K context length Python dataset (1.875M function-level code snippets from GitHub, Codeparrot, and GPT4-o-generated data) Software Security High accuracy in detecting vulnerabilities across 14 CWEs, F1 scores ranging from 84% to 99% Further improvement in identifying complex vulnerabilities and handling diverse programming patterns
Ding et al. [89] 2025 LLM-enhanced framework with in-context learning and CoT reasoning SolidiFI benchmark dataset Blockchain Security Recall of 95.06% and F1-score of 94.95% for detecting smart contract vulnerabilities; self-check architecture for CoT generation Expanding the framework’s applicability to more complex blockchain environments and new vulnerability types
Arshad et al. [90] 2025 LLM-based decentralized vehicular network architecture Simulation data for vehicular communication scenarios Autonomous Transportation Systems 18% reduction in latency, 12% improvement in throughput, and enhanced secure V2X communication using blockchain and LLMs Addressing node selfishness, scalability in larger networks, and privacy-preserving real-time data exchange
Xiao et al. [91] 2025 LLM with advanced prompting techniques Solidity v0.8 vulnerabilities dataset Blockchain Security Reduced false-positive rates by over 60%; evaluated latest five LLMs and identified root causes for reduced recall in Solidity v0.8 Improving recall for newer Solidity versions and adapting to evolving library and framework changes
Hassanin et al. [92] 2025 Pre-trained Transformer with specialized input transformation module UNSW_NB 15 and TON_IoT datasets Intrusion Detection Achieves 100% accuracy on UNSW_NB 15 dataset, significantly outperforming BiLSTM, GRU, and CNN models Exploring scalability for larger and more diverse datasets; integrating real-time detection capabilities
Liu et al. [93] 2025 LLM-powered static binary taint analysis Real-world firmware datasets Hardware Security Bugs Fully automated taint analysis with 37 newly discovered bugs and 10 assigned CVEs; low engineering cost Exploring adaptability for diverse binary formats and enhancing real-time analysis capabilities
Gaber et al. et al. [94] 2025 Transformer-based framework for zero-day ransomware detection Assembly instructions captured by the Peekaboo tool Malware Classification Introduced a novel AI-based framework leveraging Assembly data for high-accuracy zero-day ransomware detection; demonstrated the relevance of Transformer models to ransomware classification by aligning with Zipf’s law Enhancing scalability for larger datasets and addressing advanced evasion techniques in novel ransomware samples
@@ IV LLMs-based models for Cyber Security
This section reviews recent studies employing LLM-based models (i.e., Recurrent Neural Networks-based and transformer-based models) for threat detection, malware classification, intrusion detection, and software vulnerability detection.Table II presents the RNN-based models for Cyber Security, while Tables III and IV present the transformer-based models for Cyber Security. Figure 4 presents the LLM-based solutions for Cyber Security Use Cases.
@@ IV-B1 Cloud Threat Forensics
Karlsen et al. [87] proposed the LLM4Sec framework that demonstrates the potential of large language models in cybersecurity log analysis by benchmarking 60 fine-tuned models, including architectures like BERT, RoBERTa, GPT-2, and GPT-Neo. The study highlights the importance of fine-tuning for domain adaptation, with DistilRoBERTa achieving an exceptional F1-score of 0.998 across diverse datasets. This work introduces a novel experimentation pipeline that can serve as a foundation for further advancements in automated log analysis. Future research could focus on scaling these models to handle various log formats and optimizing them for real-time, dynamic cybersecurity environments.
@@ IV-B2 Malware classification
Patsakis et al. [84] investigates the application of LLMs in malware deobfuscation, focusing on real-world scripts from the Emotet malware campaign. The evaluation highlights the potential of LLMs in identifying key indicators of compromise, achieving 69.56% accuracy for URLs and 88.78% for associated domains. These findings emphasize the importance of fine-tuning LLMs for specialized cybersecurity tasks, such as reverse engineering and malware analysis. While promising, the work identifies areas for improvement, including optimizing fine-tuning strategies to enhance accuracy and integrating these capabilities into threat intelligence frameworks for real-world application.
Gaber et al. et al. [94] proposed a Pulse framework that pioneers the use of Transformer models for zero-day ransomware detection by analyzing Assembly instructions captured through the Peekaboo dynamic binary instrumentation tool. By leveraging Zipf’s law, the study effectively connects linguistic principles with ransomware behavior, making Transformer models ideal for classification tasks. This innovative approach forces the model to focus on malicious patterns by excluding familiar functionality, ensuring robust detection of novel ransomware. Future research could expand scalability to accommodate larger datasets and address increasingly sophisticated evasion techniques in emerging ransomware threats.
@@ IV-B3 Intrusion Detection
Wu et al. [68] proposed an RTID that reconstructs feature representations in imbalanced datasets to make a trade-off between dimensionality reduction and feature retention. The proposed method utilizes a stacked encoder-decoder neural network and a self-attention mechanism for network traffic type classification. The results with CICID2017 and CIC-DDoS2019 datasets demonstrate the proposed method’s effectiveness in intrusion detection compared to classical machine learning and deep learning algorithms. Ghourbi et al. [70] propose an intrusion and malware detection system to secure the entire network of the healthcare system independently of the installed devices and computers. The proposed solution includes two components: an intrusion detection system for medical devices installed in the healthcare network and a malware detection system for data servers and medical staff computers. The proposed system is based on optimized LightGBM and Transformer-based models. It is trained with four different datasets to ensure a varied knowledge of the different attacks affecting the healthcare sector. The experimental evaluation of the approach showed remarkable accuracies of 99%. PLLM-CS [92] introduces a transformative approach to satellite network security, achieving perfect accuracy on a benchmark dataset and demonstrating superior performance over traditional deep learning models.
@@ IV-B4 Software Vulnerability Detection
SecureQwen [88] is a vulnerability detection system designed for Python codebases. It uses a decoder-only transformer model with an extended context length of 64K tokens to analyze large-scale datasets. The model identifies vulnerabilities across 14 types of CWEs with high accuracy, achieving F1 scores ranging from 84% to 99%. By leveraging a dataset of 1.875 million function-level code snippets from various sources, including GitHub and synthetic data, SecureQwen demonstrates its capability to detect security issues in both human-written and AI-generated code.
Guo et al. [85] explores the role of LLMs in detecting vulnerabilities in source code, comparing the performance of fine-tuned open-source models and general-purpose LLMs. Leveraging a binary classification task and multiple datasets demonstrates the importance of fine-tuning smaller models for specific tasks, sometimes outperforming larger counterparts. The analysis also exposes critical issues with current benchmark datasets, such as mislabeling, which significantly affects model training and evaluation. Future research directions include improving dataset quality and developing strategies to enhance model generalization for more diverse and complex software vulnerabilities.
Lykousas and Patsakis [86] examine developer password patterns and the role of LLMs in detecting hard-coded credentials in source code. The study reveals that while developers tend to select more complex passwords compared to regular users, context often influences weaker patterns. It underscores the risks posed by public repositories containing secrets and the need for enhanced security practices. Additionally, the paper evaluates LLMs for detecting hard-coded credentials and identifying their potential and limitations. Future work should focus on refining LLM capabilities to detect sensitive information and raising developers’ awareness about secure password management.
@@ IV-B7 Hardware Security Evaluation
Madhav et al. [101] evaluated the security of the HDL code generated by ChatGPT. The authors introduced a similar taxonomy to the NIST CWE  [^1] https://nvd.nist.gov/vuln/categories. The authors conducted various experiments to explore the impact of prompt engineering on the security of the generated hardware design.
Liu et al. [93] introduces a groundbreaking approach, named LATTE, to binary program security by utilizing LLMs for static binary taint analysis. Unlike traditional tools like Emtaint and Karonte, which rely on manually crafted taint propagation and vulnerability inspection rules, LATTE is fully automated, reducing dependency on human expertise. Its effectiveness is demonstrated through the discovery of 37 previously unknown bugs in real-world firmware, with 10 earning CVE assignments. Additionally, LATTE offers a scalable and cost-efficient solution, making it highly accessible to researchers and practitioners. This work highlights the potential of LLMs to revolutionize binary program analysis, though future research could focus on enhancing adaptability to diverse binary formats and integrating real-time capabilities.
@@ IV-B9 Protocols verification
IV-B10 Blockchain Security
SmartGuard [89] is a framework that combines large language models with advanced reasoning techniques to detect vulnerabilities in smart contracts. It uses semantic similarity to retrieve relevant code snippets and employs Chains of Thought (CoT) reasoning for in-context learning. The framework includes a self-check mechanism for generating reliable reasoning chains from labeled data. Tests on the SolidiFI benchmark dataset show exceptional results, with a recall of 95.06% and an F1-score of 94.95%, outperforming existing tools in smart contract security. BlockLLM [90] introduces a decentralized network architecture for autonomous vehicles, integrating blockchain with large language models to improve security and communication. It enhances vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) communication by providing adaptive decision-making and ensuring data integrity. With features like incentive mechanisms for node reliability, BlockLLM achieves significant improvements, including an 18% reduction in latency and a 12% increase in throughput, offering a scalable solution for secure vehicular networks. Xiao et al. [91] advances the field of smart contract vulnerability detection by focusing on Solidity v0.8, the latest version, unlike earlier works based on outdated versions. By leveraging advanced prompting techniques with five cutting-edge LLMs, the study significantly reduces false-positive rates (over 60%), showcasing the potential of refined LLM utilization. However, the findings also reveal a significant drop in recall for specific vulnerabilities due to challenges adapting to newly introduced libraries and frameworks. Addressing these limitations could further enhance the precision and robustness of LLM-based smart contract analysis.
TABLE V: Comparison of Large Language Models
TABLE VI: Continued
TABLE VII: Continued
V General LLMs
Tables V, VI, VII compare general transformer-based Large Language Models. LLM models are generally trained on a diverse and broad range of data to provide a relatively comprehensive understanding. They can handle various language tasks like translation, summarization, and question-answering. In contrast, code-specific LLMs are specialized models trained primarily on programming languages and related technical literature, which makes their primary role in understanding and generating programming code well-suited for tasks like automated code generation, code completion, and bug detection.
V-A Prevalent LLMs
V-A1 GPT-3
V-A2 GPT-4
V-A3 T5
V-A4 BERT
V-A5 ALBERT
V-A6 RoBERTa
V-A7 XLNet
V-A8 ProphetNet
V-A9 Falcon
V-A10 Reformer
V-A11 PaLM
V-A12 PaLM2
V-A13 LLaMA
V-A14 LLaMA2
V-A15 GShard
V-A16 ELECTRA
V-A17 MPT-30B
V-A18 Yi-34B
V-A19 Falcon2-11B
TABLE VIII: Comparison of 19 LLMs Models’ Performance in Hardware Security Knowledge.
@@ V-A9 Falcon
V-B LLMs performance in the hardware cybersecurity
Table VIII compares 19 publicly available LLMs’ performance in Hardware design-related bug detection and security issues identification using samples from various sources. A portion of the Chrystalis dataset [102] has been used to evaluate the performance of the LLM models in bug detection tasks. A set of faults has been injected intentionally into a functional code and labeled as faulty. The sample size that has been processed comprises 10K of hardware design-related code samples. The prompt that has been used instructs the model to check the concerned code for any issue or bug and respond only with yes or no. The result is presented as the ratio of the responses where the model successfully identified a buggy code from the total samples used. The hardware CWE column evaluates the capability of the models to link a buggy code to its CWE number. The prompt has been designed to ask for a well-defined CWE number on the buggy design. This evaluation process asses the capability of an LLM model in bug detection and classification into the correct CWE class number.
TABLE IX: Comparison of 42 LLMs Models’ Performance in Cyber Security Knowledge.